You earn $210 in 40 hours at this rate how much do you earn in 55 hours?

Answer by oberobic(2304)

$210 for 40 hours = 210/40 = 21/4 = 5 1/4 = $5.25 per hr.
.
21/4 * 55 = 5.25*55 = $288.75
.
You can check this multiplying 15*5.25 and adding to $210.
$78.75 + $210 = $288.75.
Correct.